# Over-The-Air Double-Threshold Deep Learner for Jamming Detection in 5G RF domain

### 05 Mar 2024 v1

@@ V-A3 DT-DDNN Performance
The confusion matrix of the DT-DDNN design which includes a double DNN architecture to improve the sensitivity of the jamming detector is presented in Fig. 12. While the first DNN uses double-threshold concept to increase the sensitivity, the second DNN includes a deep cascade learning model to enhance the ability of the system to accurately classify complex observations and decrease the probability of miss-detection to 9.0%. The model accurately classifies 91.0% of non-jammed cases and 96.4% of jammed observations, which shows almost 3% and 10.4% rise in the accuracy compared to the single threshold architecture. These results represent an advancement in the ability to detect and provide countermeasures for jamming attacks in real-world scenarios of the 5G network.
Comparative analysis of three DNN designs for jamming detection in a 5G network is provided by the ROC curves in Fig. 13. The green dotted curve corresponds to a DNN trained on raw IQ samples taken from 5G waveform with no preprocessing block in the design. The optimum configuration achieved with 4 layers of CNN with batch and ReLu following each layer. The single threshold DNN design, illustrated by the blue dashed curve, is correspondence to the single threshold DNN design that uses a single DNN to classify jammed cases. In contrast, the red solid curve represents the DT-DDNN design which uses a double DNN system to improve the sensitivity of the detection algorithm. The primary DNN block detects jamming be applying a double threshold decision-making method, and the secondary DNN handles observations which fall into the ambiguous area that the first DNN has found difficulties in classifying. By integrating a deep cascade learning model into the second DNN, the objective is to further enhance the classification performance. Based on the ROC curves, it is evident that the DT-DDNN provides a higher probability of detection () in comparison to the majority of false alarm probabilities (). This indicates a greater proportion of true positive and a reduced potential of miss-detection of legit signals as jamming. This suggests that the second design with DT-DDNN architecture offers a more robust jamming detection system under challenging detection cases.
@@ V-B Experimental Validation
To extract precise information of SSB, it is critical to perform time offset (TO) and carrier frequency offset (CFO) estimations. This is due to the lack of knowledge on exact center frequency, which necessitates the application of a blind search. To accurately calculate the TO and CFO, we utilize the PSS correlation characteristics along with the cyclic prefix from CP-OFDM 5G waveform to fine tune with gNB signal. The optimization problem for estimating CFO is formulated as,
@@
respectively. Following the extraction of SSB, the correlation signals and EPNRE values are calculated. Fig. 15 demonstrates OFDM symbols of SSB versus the subcarriers indices for one of the observations selected as an example. In this figure, black cross markers represent the PSS symbols which are used to calculate PSS correlation, and red dots are correspondent to null subcarriers which provide EPNRE values. The features of collected samples, including 6000 observations, are then transferred into a 3D tensor and fed into the DT-DDNN model. The threshold parameters of the DT-DDNN which defines the sensitivity of the model are updated based on environmental noise power.
The confusion matrix in Fig. 16 provides an evaluation of the performance of DT-DDNN model using data obtained from the practical testbed. Based on these results, the model accurately classifies 93.6% of non-jammed cases and 94.1% of jammed observations. 213 observation of jammed signals are classified as non-jammed cases with miss-detection rate of 6.4% and 217 cases of non-jammed observation are classified as jamming observation with false alarm probability of 5.9%. These results validate the practical applicability and efficacy of DT-DDNN model by showing the ability of the proposed model to accurately distinguish between jammed and non-jammed 5G signal in experimental configuration.

### 11 Mar 2024 v2

@@ V-A3 DT-DDNN Performance
The confusion matrix of the DT-DDNN design which includes a double DNN architecture to improve the sensitivity of the jamming detector is presented in Fig. 12. While the first DNN uses a double-threshold concept to increase the sensitivity, the second DNN includes a deep cascade learning model to enhance the ability of the system to accurately classify complex observations and decrease the probability of miss-detection to 9.0%. The model accurately classifies 91.0% of non-jammed cases and 96.4% of jammed observations, which shows almost 3% and 10.4% rise in the accuracy compared to the single threshold architecture. These results represent an advancement in the ability to detect and provide countermeasures for jamming attacks in real-world scenarios of the 5G network.
Comparative analysis of three DNN designs for jamming detection in a 5G network is provided by the ROC curves in Fig. 13. The green dotted curve corresponds to a DNN trained on raw IQ samples taken from 5G waveform with no preprocessing block in the design. The optimum configuration was achieved with 4 layers of CNN with batch and ReLu following each layer. The single threshold DNN design, illustrated by the blue dashed curve, is correspondence to the single threshold DNN design that uses a single DNN to classify jammed cases. In contrast, the red solid curve represents the DT-DDNN design which uses a double DNN system to improve the sensitivity of the detection algorithm. The primary DNN block detects jamming by applying a double threshold decision-making method, and the secondary DNN handles observations that fall into the ambiguous area that the first DNN has found difficulties in classifying. By integrating a deep cascade learning model into the second DNN, the objective is to further enhance the classification performance. Based on the ROC curves, it is evident that the DT-DDNN provides a higher probability of detection () in comparison to the majority of false alarm probabilities (). This indicates a greater proportion of true positive and a reduced potential of miss-detection of legit signals as jamming. This suggests that the second design with DT-DDNN architecture offers a more robust jamming detection system under challenging detection cases.
@@ V-B Experimental Validation
To extract precise information of SSB, it is critical to perform time offset (TO) and carrier frequency offset (CFO) estimations. This is due to the lack of knowledge on the exact center frequency, which necessitates the application of a blind search. To accurately calculate the TO and CFO, we utilize the PSS correlation characteristics along with the cyclic prefix from CP-OFDM 5G waveform to fine-tune with gNB signal. The optimization problem for estimating CFO is formulated as,
@@
respectively. Following the extraction of SSB, the correlation signals and EPNRE values are calculated. Fig. 15 demonstrates OFDM symbols of SSB versus the subcarrier indices for one of the observations selected as an example. In this figure, black cross markers represent the PSS symbols that are used to calculate PSS correlation, and red dots correspond to null subcarriers from which EPNRE is calculated. The features of collected samples, including 6000 observations, are then transferred into a 3D tensor and fed into the DT-DDNN model. The threshold parameters of the DT-DDNN which defines the sensitivity of the model are updated based on environmental noise power.
The confusion matrix in Fig. 16 provides an evaluation of the performance of DT-DDNN model using data obtained from the practical testbed. Based on these results, the model accurately classifies 93.6% of non-jammed cases and 94.1% of jammed observations. 213 observations of jammed signals are classified as non-jammed cases with a miss-detection rate of 6.4% and 217 cases of non-jammed observation are classified as jammed observations with false-alarm probability of 5.9%. These results validate the practical applicability and efficacy of DT-DDNN model by showing the ability of the proposed model to accurately distinguish between jammed and non-jammed 5G signals in experimental configuration.